OpenStudy (anonymous):

Can someone please explain to me how to find the exact values of sin 7/4 pi. I'm so confused with converting radian measure to degree measure.

OpenStudy (anonymous):

pi = 180

OpenStudy (anonymous):

ohh.. it's not 3.14? o.O Why is pi = 180? and not 3.14

OpenStudy (anonymous):

What you need to know when converting is 2pi=360degrees. http://www.mathwarehouse.com/trigonometry/radians/convert-degee-to-radians.php That site explains it better than I can.

OpenStudy (anonymous):

\[\frac{\pi}{4}=45^o\] and \[\frac{7\pi}{4}=-\frac{\pi}{4}\] sin in 4th quadrant is negative so what is \[\sin(\frac{\pi}{4})=?\]